OpenStudy (anonymous):

Write the following quadratic function in vertex form. Determine the vertex and the axis of symmetry for the function f(x)=-5x^2+7x+9

OpenStudy (tkhunny):

Have you considered "Completing the Square"?

OpenStudy (anonymous):

Yes, I always get messed up and I just can't complete the square. It's been something I've been trying to work on..

OpenStudy (tkhunny):

Well, let's walk through it. First, you have to get this burned into your brain: \((a+b)^{2} = a^{2} + 2ab + b^{2}\) Stare at the 'b'-parts.

OpenStudy (anonymous):

Okay, now what?

OpenStudy (tkhunny):

There are a few ways to go about it, bu I like to make the leading coefficient one (1). f(x)=-5x^2+7x+9 = -5(x^2 - (7/5)x) + 9 That's pretty ugly, but now we just need to work inside the parentheses.

OpenStudy (anonymous):

Yeah that looks exactly how my teacher would do it. I just could never understand what to do.

OpenStudy (anonymous):

use the formula for vertex form of y=a(x-h)^2+k, a=-5 , b=7 c=9, vertex x=-b/2a =7/10 yes or no?whats k or y coordinate of vertex?

OpenStudy (anonymous):

your vertex in that formula is V(h,k)

OpenStudy (tkhunny):

Now we need that thing you have burned into your brain. Let's hold them up next to each other and see what it looks like. \(a^{2} + 2ab + b^{2}\) \(x^{2} - \frac{7}{5}x\) Remember those b-parts? The \(b^{2}\) is missing. The plan is to use the 'ab' term to build it.

OpenStudy (anonymous):

mark_o I'm not really understanding how you are solving it. Thank you for your help but I think I'm going to stick with the way that tkhunny is walking me through it since it is more like the way my teacher does it.

OpenStudy (anonymous):

How do you build it? this is the part I get stuck at the most...

OpenStudy (anonymous):

if vertex x=7/10 then sub this to f(x)=y=-5x^2+7x+9, is y=k=229/20=11.45 yes or no ? then sub these h and k to the formula of vertex form y=a(x-h)^2+k

OpenStudy (anonymous):

a=-5,h=7/10=0.7, k=229/20=11.45 just plug them to the formula y=a(x-h)^2+k

OpenStudy (tkhunny):

See how the \(x^{2}\) matches up with the \(a^{2}\)? This gives us permission to pull the 'x' out of the middle term and be left with the '2b' part, So, \(x^{2}\) relates to \(a^{2}\), And (-7/5)x relates to 2ab. This says 2b = -7/5.

OpenStudy (anonymous):

okay but how do i find the a.o.s and the vertex?

OpenStudy (tkhunny):

Let's not get ahead of ourselves. We still need to find the b^2 -part. We have 2b = -7/5. Well, then b = -7/10. This leads immediately to b^2 = 49/100.